OpenStudy (anonymous):

Find the sum of infinite series : 1+(1+b)r+(1+b+b^2)r^2 + (1+b+b^2+b^3)r^3+...... where r and b are proper fractions.

OpenStudy (ash2326):

Let the sum be S \[S=1+ (1+b)r+(1+b+b^2)r^2+(1+b+b^2+b^3)r^3+...\] multiply S by r, we get \[rS=r+(1+b)r^2+(1+b+b^2)r^2+(1+b+b^2)r^3+....\] now subtract S by rS we get \[S-rS=1+br+b^2r^2+b^3r^3+...\] or \[S(1-r)=1+br+b^2r^2+b^3r^3\] could you solve for s now, the right side is just GP

OpenStudy (ash2326):

@shubham.bagrecha do you get this?

OpenStudy (anonymous):

yes